OpenStudy (jkbo):

Find the distance between points M(6, 16) and Z(–1, 14) to the nearest tenth.

OpenStudy (anonymous):

\[\sqrt{(x _{1} - x _{2})^{2} + (y _{1} - y _{2})^{2}}\] \[\sqrt{(6 - (-1))^{2} + (16 - 14)^{2}}\] \[\sqrt{(7)^{2} + (2)^{2}}\] \[\sqrt{49 + 4}\] \[\sqrt{53}\]